OpenStudy (anonymous):

5 x squared over the difference of x and 4 · the difference of 2 x squared and 32 all over 20 x to the fourth PLEASE HELP, I NEED IT!! PLEASE EXPLAIN!! :-)

OpenStudy (anonymous):

20 a cubed b plus 15 a b to the fourth minus 25 a b all over 5 a b?

OpenStudy (mertsj):

\[\frac{20a^3b+15ab^4-25ab}{5ab}\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

Yes it is :-)

OpenStudy (mertsj):

\[\frac{20a^3b}{5ab}+\frac{15ab^4}{5ab}-\frac{25ab}{5ab}\]

OpenStudy (mertsj):

\[4a^2+3b^3-5\]

OpenStudy (anonymous):

ty on his behalf, lol lol lol